how would I Find the equation (in terms of x) of the line through the points
(-4,-4) and (2,5)

Answer by god2012(113)

Given:(x1,y1) = (-4,-4) and (x2,y2) = (2,5)
2-point form of the equation of a line is :
(y2-y1)/(x2-x1) = (y-y1)/(x-x1)
(5-(-4))/(2-(-4)) = (y+4)/(x+4) = 9/6 = 3/2
2(y+4) = 3(x + 4)
3x - 2y - 4 = 0 is the equation of the line